i am a five digit number. my ones digit is 7.my thousand digit is 5 less than the ones digit.
my tens digit is 2more than the thousand digit.
my ten thousand digit is 1 more than the thousand digit
my hundreds digit is 6 more than the trns digit

Answer by Edwin McCravy(20077)

>>i am a five digit number.<< 

So we put 5 blanks for the digits.

_ _ _ _ _

The first one is the ten thousands digit.
The second one is the thousands digit.
The third one is the hundreds digit.
The fourth one is the tens digit.
The fifth one is the ones digit.
_ _ _ _ _

>>my ones digit is 7.<< 

So we put a 7 for the fifth digit:

_ _ _ _ 7

>>my thousand digit is 5 less than the ones digit.<<

The ones digit is 7.  To get a number which is 5 less,
we subtract 5 from 7.  7 - 5 = 2.  So we put 2 for the 
second digit:

_ 2 _ _ 7

>>my tens digit is 2 more than the thousand digit.<<

The thousands digit is 2 and two more than 2 is found by
adding 2+2 = 4, so the tens digit is 4 and we put 4
for the fourth digit:

_ 2 _ 4 7

>>my ten thousand digit is 1 more than the thousand digit<<

The thousands digit is 2 and 1 more that 2 is found by adding
2+1 = 3 so the ten thousands digit is 3. So we put the first
dogit as 3:

3 2 _ 4 7

>>my hundreds digit is 6 more than the trns digit<<

I don't know what you meant by "trns" but it can't be "tens" because
the tens digit is 4 and 6 more than 4 is 4+6=10, which is not a digit.
"trns" can't mean "ones" because the ones digit is 7 and 7+6=13, which
is not a digit.

So we are stuck here until we find out what "trns" is supposed to be.
If it is "tens" or "ones" then there is no possible answer.  It would 
only make sense if "trns" means either "thousands" or "ten thousands".

If "trns" means "thousands", then 6 more than 2 is found by adding 2+6=8,
so in that case the answer is

3 2 8 4 7

If "trns" means "ten thousands", then 6 more than 3 is found by adding 3+6=9,
so in that case the answer is

3 2 9 4 7

The typo spoiled the problem.  But did you really find this to be too difficult
for you?  It is very easy if you know which digit is which.

Edwin